How many grams of beryllium chloride are needed to make 125mL of a 0.050 M solution
1
Expert's answer
2013-03-29T04:41:33-0400
The equation needed to calculate theconcentration is

C(M) = n(mol) / V(L)

C - concentration (M)
n - amount of BeCl2 (mol)
V - volume (L)

n(mol) = C(M) x V(L) = 0.050 x 0.125 = 6.25 x 10^-3 mol

The mass of BeCl2 is

m(g) = n(mol) x MW(g/mol)

MW(BeCl2) = 9.01 + 2 x 35.5 = 80.01 g/mol
m(BeCl2) = 6.25 x 10^-3 x 80.01 = 0.500 g

Answer: m(BeCl2) = 0.500g